OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (babalooo):

\[\frac{ 4 }{ 5 }x + \frac{ 4 }{ 3 } = 2x\] Isolate variable terms to one side such that 4/3 = 2x- 4x/5 To add/subtract fractions, we need same denominator so 4/3 = 10x/5 -4x/5 4/3 = 6x/5 Now we isolate variable x by multiply both sides by 5 and divide both sides by 6 4/3 * 5/6 = 6x/5 *5/6 20/18 = x Now simplify 20/18 and that's the value of x